OVHCloud Bare Metal Cloud Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5879 — VPS: Switching the infrastructure
Scheduled Maintenance Report for Bare Metal Cloud
Completed
We will implement new N5548 to replace N5010.Then, we will switch the FEX of the rack on the new N5. During this time, traffic will switch to the other side.

Before that, we have to change the configuration of routers.

Update(s):

Date: 2011-10-19 10:11:18 UTC
During the maintenance (since 1 day), we lost
14 packets

--- 46.105.24.104 ping statistics ---
91579 packets transmitted, 91565 received, 0% packet loss, time 91895190ms

Except for the case of host which was poorly redundant:

--- 46.105.7.106 ping statistics ---
91681 packets transmitted, 90207 received, +2 errors, 1% packet loss, time 91993321ms

24 minutes of crashes. it is a VPS client who took the
99.9% offer and did not (yet) switched on the HA offers
ensuring 99.99% for a little extra money.

Date: 2011-10-19 10:06:14 UTC
The configuration of side B was modified. the router was re-put
on the routing. Both of the routers are active
both of the A and B infra are active.

End of the maintenance.

Date: 2011-10-07 14:57:12 UTC
We have just cut the old infrastructure.

Date: 2011-10-07 14:56:47 UTC
The traffic is cut on the s10.Everything is routed by the s9.
We are doing the configuration change.

Date: 2011-10-07 14:55:13 UTC
The B side is up. We have finalized the setting of 4x10G
per rack.

We switch the routing configuration on the rbx-s10-6k.
So we will cut traffic on the s10. It's going to work
on the s9. Then we change the configuration.

Date: 2011-10-07 14:53:44 UTC
The B side is up. We have finalized the setting of 4x10G
per rack.

We switch the routing configuration on the rbx-s10-6k.
So we will cut traffic on the s10. It's going to work
on the s9. Then we change the configuration.

Date: 2011-10-07 14:49:23 UTC
We switch on the B side.

Date: 2011-10-07 14:48:59 UTC
We have finalised the B side,it will take roughly 2h.

Date: 2011-10-07 14:47:23 UTC
Everything is cut on the B side.Every thing work on the A side.

Date: 2011-10-07 14:46:39 UTC
We have forgotten to reconfigure the router.Fixed

Date: 2011-10-07 14:28:50 UTC
The A side is connected on the vps-20-n5
We attack the B side ie we upgrade FEX of vps-11-n5 on vps-21-n5.
We cut the B port on host

Date: 2011-10-07 14:26:30 UTC
A host is not correctly connected on the B side.

Date: 2011-10-07 14:25:37 UTC
The A side is cut on the vps-10-n5.We disconnect on the vps-20-n5

Date: 2011-10-07 14:24:15 UTC
We restart the upgrade from the A side.
vps-10-n5 vers vps-20-n5

Date: 2011-10-07 05:08:51 UTC
The 2a/b are ready,the manipulation will be done tomorrow.

Date: 2011-10-07 05:08:06 UTC
Traffic is now on rbx-s10 only.We will change the configuration and connect properly the new vps-2a/b and its satellite vss-20/21.

Date: 2011-10-07 05:05:22 UTC
We will cut the traffic on rbx-s9-6k
Posted Oct 07, 2011 - 05:04 UTC